# Interactive Query Answering on Knowledge Graphs with Soft Entity Constraints

### v2

@@ Abstract
Methods for query answering over incomplete knowledge graphs retrieve entities that are likely to be answers, which is particularly useful when such answers cannot be reached by direct graph traversal due to missing edges. However, existing approaches have focused on queries formalized using first-order-logic. In practice, many real-world queries involve constraints that are inherently vague or context-dependent, such as preferences for attributes or related categories. Addressing this gap, we introduce the problem of query answering with soft constraints. We formalize the problem and introduce two efficient methods designed to adjust query answer scores by incorporating soft constraints without disrupting the original answers to a query. These methods are lightweight, requiring tuning only two parameters or a small neural network trained to capture soft constraints while maintaining the original ranking structure. To evaluate the task, we extend existing QA benchmarks by generating datasets with soft constraints. Our experiments demonstrate that our methods can capture soft constraints while maintaining robust query answering performance and adding very little overhead.
@@ 1 Introduction
To tackle these challenges, we formalize the problem and introduce efficient methods designed to capture soft constraints while preserving the global structure in an original list of scores for a given query. In summary, our work makes the following key contributions:
@@
    We extend existing benchmarks for query answering with automatically generated data of soft constraints. Through comprehensive evaluations, we show how our methods effectively capture these constraints without significantly compromising global ranking performance or runtime performance.
@@ 4 Query Answering with Soft Constraints
Definition 1 (Soft constraints (informal)).
@@
Here, the label $l=1$ marks Best Sound as a positive example (entities similar to it (e.g., Best Costume Design, Best Original Song) should receive higher scores) while $l=0$ marks Best Picture as a negative example, encouraging the model to de-emphasize similar entities (e.g., Best Director). After incorporating $P(2)$, the updated scores $a^{(2)}(q)$ should give priority to entities that reflect the soft constraint.
@@ 5 Incorporating Soft Constraints
Let $n=|\mathcal{V}|$ be the number of entities, $d$ the embedding dimension, and $t=|P(t)|$ the number of preferences. For each entity $e\in\mathcal{V}$ we compute $\delta(P^{\odot},e)$ as the mean of $t$ cosine similarities. Each cosine reduces to a dot product between $d$-dimensional vectors, so it costs $\mathcal{O}(d)$; hence one entity requires $\mathcal{O}(kd)$ time to accumulate $p$ dot products and take their mean, and processing all $n$ entities costs $\mathcal{O}(nkd)$ time.
@@
Intuitively, making the weights conditional allows to contextualize the score update to specific values of scores, similarities, and the entity as captured by its embedding. The MLP outputs two scalars $\alpha_{p},\alpha_{n}\in(0,1)$, which we then use in the update rule. We refer to this as the NQR update:
@@
If the hidden dimension of the MLP is $h$, the forward pass per entity multiplies a $(d+3)\times h$ matrix and an $h\times 2$ matrix (plus activations), which is $\mathcal{O}(dh)$ time; across all entities this adds $\mathcal{O}(ndh)$. Thus the NQR update over all entities runs in $\mathcal{O}(n(kd+dh))$ time, remaining linear in the number of entities in the KG.
@@
The main goal of NQR is to adjust entity scores such that the information conveyed by the preference data is reflected in the ranking, while preserving the overall quality of the original query answers. This can be achieved with a training set of logical queries and their corresponding answers, together with a partition of of the set of answers into disjoint preference sets $P^{+}$ and $P^{-}$. The training objective is then two-fold: (1) entities in $P^{+}$ should be ranked higher than those in $P^{-}$, and (2) all entities in $P^{+}\cup P^{-}$ (which satisfy the original logical query) should still rank above entities that do not meet the logical constraints of the query.
To achieve this, we adopt the RankNet loss (burges_learning_2005), which encourages pairs of relevant and non-relevant items to be ordered correctly. Given two entities $e_{i}$ and $e_{j}$ with scores $a[e_{i}]$ and $a[e_{j}]$, RankNet minimizes the binary cross-entropy over the probability $\sigma(a[e_{i}]-a[e_{j}])$ that $e_{i}$ should be ranked above $e_{j}$.
For NQR, we define two complementary loss terms:
@@ 6.1 Datasets
Prior work on query answering on KGs has relied on datasets containing query-answer pairs $(q,\mathcal{A})$, where $q$ is a query and $\mathcal{A}\subset\mathcal{V}$ is the set of answers computed over a KG (ren_query2box_2020; ren_beta_2020). We extend these datasets by deriving preference data from observable clusters of entities in the set $\mathcal{A}$. This process consists of three steps (we provide additional details in Appendix  B):
@@
2. Clustering. We apply Hierarchical Agglomerative Clustering using cosine similarity as the distance metric. As shown in Fig. 2, the result is a dendrogram that iteratively merges the most similar entities into clusters, capturing similarities at different levels.
3. Answer set partitioning. We traverse the dendrogram from the root node towards the leaves, and if a cluster contains at least 20% of the entities in the answer set [^1] We define this limit to avoid clusters with too few entities., we use it to form a preference set. For each entity $e_{i}\in\mathcal{A}$, we define its preference label as $l_{i}=1$ if it is in the cluster, and $l_{i}=0$ otherwise. An example of a preference set is shown in Fig. 2, with entities labeled as $l_{i}=1$ are highlighted in green, and the rest of the entities are labeled as $l_{i}=0$. This partition forms the preference set $P(T)=\{(e_{1},l_{2}),\ldots,(e_{T},l_{T})\}$. We then denote an instance in our extended datasets as $(q,\mathcal{A},P(T))$, with instances having values of $T$ between 10 and 100.
Prior work has relied on subsets of encyclopedic KGs such as NELL (carlson_toward_2010) and Freebase (bollacker_freebase_2008) to benchmark methods for query answering. These KGs contain statements about people, organizations, and locations around the world. For our experiments, we select FB15k237 (toutanova_observed_2015) and the complex queries generated for it by ren_beta_2020. To further explore the potential of query answering in a different domain we use Hetionet (himmelstein_systematic_2017), a biomedical KG containing $\sim$2M edges between genes, diseases, and drugs, among other types. For Hetionet, we extract complex queries of different types, and then for both datasets we run the three steps above to collect preference data. We consider 14 types of first order logic queries of varying complexity, including negation. We present statistics of the final datasets and splits in Table 1. For training NQR, we only use the single type of 1-hop, link prediction queries. For the validation and test sets we use the full set of 14 types of complex queries. More detailed statistics can be found in Appendix  C.
@@ 6.2 Evaluation
2. Ranking metrics: Following prior work on CQA, we evaluate evaluate query answering performance using ranking metrics. For an answer $e\in\mathcal{A}$ to a query, the reciprocal rank is $1/\text{r}(e,a^{(t)})$, and the Mean Reciprocal Rank (MRR) is the average over all queries [^2] We compute filtered ranking metrics, where other known answers to a query are removed before computing $\text{r}(e,a^{(t)})$.. Hits at $k$ (H@k) is defined by verifying if, for each answer $e$, its rank $\text{r}(e,a^{(t)})$ is less than or equal to $k$ (in which case H@k = 1, and H@k = 0 otherwise), and averaging over all queries. We compute these ranking metrics using the original and adjusted scores to determine any effect on overall query answering performance.
@@ 6.3 Baselines
As an alternative to our proposed linear score update (Equation˜8), we consider a traditional Learning-to-Rank (LTR) baseline based on LightGBM (ke2017lightgbm), a gradient boosting framework implementing the LambdaRank objective (burges_learning_2005; burges_learning_2006). For each training query, the model receives as input three scalar features for every entity $e$: the original score $a[e]$, and the average positive and negative similarities, $\delta(P^{+},e)$ and $\delta(P^{-},e)$. Graded relevance labels are assigned following standard LTR practice (burges2010ranknet): entities in $P^{+}$ receive a label of 2, entities in $P^{-}$ receive 1, and a random set of 100 negative entities is assigned 0. The model is trained to predict ranking scores that order entities according to their relative relevance by directly optimizing NDCG via the LambdaRank objective. This baseline allows us to compare our proposed linear score update based on shifts from the original score, with a nonlinear approach that predicts the score directly using a nonlinear function of relevant features.
@@ 7 Results
We present results of pairwise accuracy, MRR, and H@10, and NDCG@10 in Fig. 3. For reference, we denote the metrics of the base QA model as Unconstrained, which allows us to determine how the adjustments in scores introduced by the methods we consider affect the performance of the original scores. We include 95% confidence intervals as shaded regions around the curves; these are narrow and often indistinguishable, indicating that the observed differences are consistent across queries.
@@
The ranking metrics (MRR and H@1) behave differently in comparison with PA. We observe that when preference sets are small, with sizes between 1 and 4, score adjustments often cause a degradation in ranking quality with respect to the original scores. The fact that PA is already high at such sizes of preference sets indicates that the adjusted scores greedily place entities in $P^{+}$ above $P^{-}$, possibly by placing wrong answers between them, which harms ranking quality. Considering all methods, and especially in Hetionet, NQR causes the smallest drops in performance at small preference sets. As preference sets grow larger, the performance of all methods (with the exception of LightGBM, which struggles on FB15k237 and in H@10 on Hetionet) increases above the original. Both Cosine and NQR yield the highest ranking metrics, indicating that linear score updates are better for preserving ranking quality in comparison with the LightGBM baseline that predicts the new score directly.
@@
We present additional results in Appendix˜E where metrics are categorized according to each of the types of complex queries in the datasets. The results show that in average Cosine performs best at balancing similarity constraints with global ranking performance, followed by NQR, which for certain query types, preserves best global ranking performance.
@@
Figure˜4 illustrates how pairwise accuracy (PA) and mean reciprocal rank (MRR) evolve jointly for different values of $\alpha$ and $\beta$ in the Cosine update (Equation˜10), as the number of interaction steps increases. Across both datasets, increasing $\alpha$ (which implies placing more weight on the original scores), prevents significant drops in MRR, but also limits the maximum PA that can be reached. Smaller $\alpha$ values, on the other hand, allow stronger preference propagation, and the effect on MRR varies depending on the value of $\beta$. This behavior reflects the trade-off in the update: higher $\alpha$ yields conservative adjustments that favor stability, while lower $\alpha$ enables more aggressive score shifts guided by similarity signals.
@@
Within each $\alpha$ setting, varying $\beta$ determines the relative emphasis placed on positive versus negative preferences. Larger $\beta$ values (yellow curves) consistently increase PA with limited loss in MRR, while negative $\beta$ values (dark red) degrade both, indicating that over-penalizing negative examples distorts the global ranking. When $\alpha=0.25$ and after 10 interactions, the values closer to the Pareto front are $\beta=0.25$ and $\beta=0$ and $\beta=0.5$. Values of $\beta=0.5$ increase PA further, at the cost of lower MRR, and viceversa for $\beta=0$. This shows that $\beta$ offers an interpretable control mechanism: users prioritizing ranking fidelity can select lower $\beta$ values to protect MRR, whereas those seeking stronger preference alignment can increase $\beta$ to boost PA. This tunability makes the Cosine update practically useful for downstream applications, enabling domain experts to adjust the balance between preserving original rankings and enforcing similarity constraints according to their needs.
@@ 8 Conclusion
We introduce efficient methods for incorporating soft constraints in complex queries. Since they reuse embeddings of a base CQA model, the methods are lightweight, requiring tuning only two parameters or a small MLP. In our experiments, we find that these methods are able to adjust query scores to capture preferences about their answers, while increasing ranking performance as more exemplary preferences are provided. In practice, we observe that our methods are fast, adding only up to 1.4 ms overhead to the base QA model.

### v3

@@ Abstract
Methods for query answering over incomplete knowledge graphs retrieve entities that are likely to be answers, which is particularly useful when such answers cannot be reached by direct graph traversal due to missing edges. However, existing approaches have focused on queries formalized using first-order-logic. In practice, many real-world queries involve constraints that are inherently vague or context-dependent, such as preferences for attributes or related categories. Addressing this gap, we introduce the problem of query answering with soft constraints. We formalize the problem and introduce two efficient methods designed to adjust query answer scores by incorporating soft constraints without disrupting the original answers to a query. These methods are lightweight, requiring tuning only two parameters or a small neural network trained to capture soft constraints while maintaining the original ranking structure. To evaluate the task, we extend existing QA benchmarks by generating datasets with soft constraints. Our experiments demonstrate that our methods can capture soft constraints while maintaining robust query answering performance and adding very little overhead. With our work, we explore a new and flexible way to interact with graph databases that allows users to specify their preferences by providing examples interactively.
@@ 1 Introduction
To tackle these challenges, we formalize the problem and introduce efficient methods designed to capture soft constraints while preserving the global structure of the original list of scores for a given query. In summary, our work makes the following key contributions:
@@
    We extend existing benchmarks for query answering with automatically generated soft constraints comprising a large and diverse set of queries with different notions of similarity. Through comprehensive evaluations, we show how our methods effectively capture these constraints without significantly compromising global ranking performance or runtime performance.
Our results point toward a more flexible interaction paradigm that complements existing methods for query answering on graphs, where answers can be adjusted according to preferences specified interactively.
@@ 4 Query Answering with Soft Constraints
Definition 1 (Soft constraint (informal)).
@@
Here, the label $l=1$ marks Best Sound as a positive example (entities similar to it, such as Best Original Song, should receive higher scores) while $l=0$ marks Best Picture as a negative example, encouraging the model to de-emphasize similar entities (e.g., Best Director). After incorporating $P(2)$, the updated scores $a^{(2)}(q)$ should give priority to entities that reflect the soft constraint.
@@ 5 Incorporating Soft Constraints
Let $n=|\mathcal{V}|$ be the number of entities, $d$ the embedding dimension, and $t=|P(t)|$ the number of preferences. For each entity $e\in\mathcal{V}$ we compute $\delta(P^{\odot},e)$ as the mean of $t$ cosine similarities. Each cosine reduces to a dot product between $d$-dimensional vectors, so it costs $\mathcal{O}(d)$; hence one entity requires $\mathcal{O}(td)$ time to accumulate $t$ dot products and take their mean, and processing all $n$ entities costs $\mathcal{O}(ntd)$ time.
@@
Intuitively, making the weights conditional allows to contextualize the score update to specific values of scores, similarities, and the entity as captured by its embedding. The MLP outputs two scalars $\alpha_{p},\alpha_{n}\in(0,1)$, which we then use in the update rule. We refer to this as the Neural Query Reranker (NQR) update:
@@
If the hidden dimension of the MLP is $h$, the forward pass per entity multiplies a $(d+3)\times h$ matrix and an $h\times 2$ matrix (plus activations), which is $\mathcal{O}(dh)$ time; across all entities this adds $\mathcal{O}(ndh)$. Thus the NQR update over all entities runs in $\mathcal{O}(n(td+dh))$ time, remaining linear in the number of entities in the KG.
@@
The main goal of NQR is to adjust entity scores such that the information conveyed by the preference data is reflected in the ranking, while preserving the overall quality of the original query answers. This can be achieved with a training set of logical queries and their corresponding answers, together with a partition of the set of answers into disjoint preference sets $P^{+}$ and $P^{-}$. The training objective is then two-fold: (1) entities in $P^{+}$ should be ranked higher than those in $P^{-}$, and (2) all entities in $P^{+}\cup P^{-}$ (which satisfy the original logical query) should still rank above entities that do not meet the logical constraints.
To achieve this, we adopt the RankNet loss (burges_learning_2005), which encourages pairs of relevant and non-relevant items to be ordered correctly. Given two entities $e_{i}$ and $e_{j}$ with scores $a[e_{i}]$ and $a[e_{j}]$, RankNet minimizes the binary cross-entropy over the probability $\sigma(a[e_{i}]-a[e_{j}])$ that $e_{i}$ should be ranked above $e_{j}$. For NQR, we define two complementary loss terms:
@@ 6.1 Datasets
Prior work on query answering on KGs has relied on datasets containing query-answer pairs $(q,\mathcal{A})$, where $q$ is a query and $\mathcal{A}\subset\mathcal{V}$ is the set of answers computed over a KG (ren_query2box_2020; ren_beta_2020). We extend these datasets by deriving preference data from observable clusters of entities in the set $\mathcal{A}$. Motivated by prior work showing that textual embeddings correlate with human notions of similarity and relatedness (2019reimerssbert; grand2022semantic; merk2022seeing), we achieve this by clustering embeddings of answers to a query, which we use as a computational proxy of similarity that allows us to study the problem at a scale of thousands of queries and comprising different notions of similarity. Concretely, this process consists of three steps (we provide additional details in Appendix  B):
@@
2. Clustering. We apply Hierarchical Agglomerative Clustering to the entities in $\mathcal{A}$, using cosine similarity as the distance metric. As shown in Fig. 2, the result is a dendrogram that iteratively merges the most similar entities into clusters, capturing similarities at different levels.
3. Answer set partitioning. We traverse the dendrogram from the root node towards the leaves, and if a cluster contains at least 20% of the entities in the answer set [^1] We define this limit to avoid clusters with too few entities., we use it to form a preference set. For each entity $e_{i}\in\mathcal{A}$, we define its preference label as $l_{i}=1$ if it is in the cluster, and $l_{i}=0$ otherwise. An example of a preference set is shown in Fig. 2, with entities labeled as $l_{i}=1$ highlighted in green, and the rest of the entities labeled as $l_{i}=0$. This partition forms the preference set $P(T)=\{(e_{1},l_{1}),\ldots,(e_{T},l_{T})\}$. We then denote an instance in our extended dataset as $(q,\mathcal{A},P(T))$, with instances having values of $T$ between 10 and 100.
Prior work has relied on subsets of encyclopedic KGs such as NELL (carlson_toward_2010) and Freebase (bollacker_freebase_2008) to benchmark methods for query answering. These KGs contain statements about people, organizations, and locations around the world. For our experiments, we select FB15k237 (toutanova_observed_2015) and the complex queries generated for it by ren_beta_2020. To further explore the potential of query answering in a different domain we use Hetionet (himmelstein_systematic_2017), a biomedical KG containing $\sim$2M edges between genes, diseases, and drugs, among other types. For Hetionet, we extract complex queries of different types, and then for both datasets we run the three steps above to collect preference data. We consider 14 types of first order logic queries of varying complexity, including negation. We present statistics of the final datasets and splits in Table 1. For training NQR, we only use 1-hop, link prediction queries. For the validation and test sets we use the full set of 14 types of complex queries. More detailed statistics can be found in Appendix  C.
@@ 6.2 Evaluation
2. Ranking metrics: Following prior work on CQA, we evaluate evaluate query answering performance using ranking metrics. For an answer $e\in\mathcal{A}$ to a query, the reciprocal rank is $1/\texttt{rank}(e,a^{(t)})$, and the Mean Reciprocal Rank (MRR) is the average over all queries [^2] We compute filtered ranking metrics, where other known answers to a query are removed before computing $\texttt{rank}(e,a^{(t)})$.. Hits at $k$ (H@k) is defined by verifying if, for each answer $e$, its rank $\texttt{rank}(e,a^{(t)})$ is less than or equal to $k$ (in which case H@k = 1, and H@k = 0 otherwise), and averaging over all queries. We compute these ranking metrics using the original and adjusted scores to determine any effect on overall query answering performance.
@@ 6.3 Baselines
We consider a traditional Learning-to-Rank (LTR) baseline based on LightGBM (ke2017lightgbm), a gradient boosting framework implementing the LambdaRank objective (burges_learning_2005; burges_learning_2006). For each training query, the model receives as input three scalar features for every entity $e$: the original score $a[e]$, and the average positive and negative similarities, $\delta(P^{+},e)$ and $\delta(P^{-},e)$. Graded relevance labels are assigned following standard LTR practice (burges2010ranknet): entities in $P^{+}$ receive a label of 2, entities in $P^{-}$ receive 1, and a random set of 100 negative entities is assigned 0. The model is trained to predict ranking scores that order entities according to their relative relevance by directly optimizing NDCG via the LambdaRank objective. This baseline allows us to compare our proposed linear score update based on shifts from the original score, with an approach that predicts the score using a nonlinear function of relevant features.
@@ 7 Results
We present results of pairwise accuracy, MRR, and H@10, and NDCG@10 in Fig. 3. For reference, we denote the metrics of the base QA model as Unconstrained, which allows us to determine how the adjustments in scores introduced by the methods we consider affect the performance of the original scores. We include 95% confidence intervals as shaded regions around the curves; these are narrow and often indistinguishable, indicating that the differences are consistent across queries.
@@
Ranking metrics (MRR and H@10) behave differently in comparison with PA. When preference sets are small, between 1 and 4 in size, score adjustments often cause a degradation in ranking quality with respect to the original scores. The fact that PA is already high at such sizes of preference sets indicates that the adjusted scores greedily place entities in $P^{+}$ above $P^{-}$, possibly by placing wrong answers between them, which harms ranking quality. Considering all methods, and especially in Hetionet, NQR causes the smallest drops in performance at small preference sets. As preference sets grow larger, the performance of all methods (with the exception of LightGBM, which struggles on FB15k237 and in H@10 on Hetionet) increases above the original. Both Cosine and NQR yield the highest ranking metrics, indicating that linear score updates are better for preserving ranking quality in comparison with the LightGBM baseline that predicts the new score directly.
@@
We present additional results in Appendix˜E where metrics are categorized according to each of the types of complex queries in the datasets. The results show that on average Cosine performs best at balancing similarity constraints with global ranking performance, followed by NQR, which for certain query types, preserves best global ranking performance.
@@
Figure˜4 illustrates how pairwise accuracy (PA) and mean reciprocal rank (MRR) evolve jointly for different values of $\alpha$ and $\beta$ in the Cosine update (Equation˜10), as the number of interaction steps increases. Across both datasets, increasing $\alpha$ (which implies placing more weight on the original scores), prevents significant drops in MRR, but also limits the maximum PA that can be reached. Smaller $\alpha$ values, on the other hand, allow stronger preference propagation, and the effect on MRR varies depending on the value of $\beta$. This behavior reflects an important trade-off: higher $\alpha$ yields conservative and stable adjustments, while lower $\alpha$ enables more aggressive shifts guided by similarity signals.
@@
Within each $\alpha$ setting, varying $\beta$ determines the relative emphasis placed on positive versus negative preferences. Larger $\beta$ values (yellow curves) consistently increase PA with limited loss in MRR, while negative $\beta$ values (dark red) degrade both, indicating that over-penalizing negative examples distorts the global ranking. When $\alpha=0.25$ and after 10 interactions, the values closer to the Pareto front are $\beta=0$ and $\beta=0.5$. Values of $\beta=0.5$ increase PA further, at the cost of lower MRR, and viceversa for $\beta=0$. This shows that $\beta$ offers an interpretable control mechanism: users prioritizing ranking fidelity can select lower $\beta$ values to protect MRR, whereas those seeking stronger preference alignment can increase $\beta$ to boost PA. This tunability makes the Cosine update practically useful for downstream applications, enabling domain experts to adjust the balance between preserving original rankings and enforcing similarity constraints according to their needs.
@@ 8 Conclusion
We introduce efficient methods for incorporating soft constraints in complex queries. Since they reuse embeddings of a base CQA model, the methods are lightweight, requiring tuning only two parameters or a small MLP. To evaluate these models over a large and diverse set of queries and notions of similarity, we extend existing query answering benchmarks with preference sets derived from clusterings of embeddings. In our experiments, we find that these methods are able to adjust query scores to capture preferences about their answers, while increasing ranking performance as more exemplary preferences are provided. In practice, we observe that our methods are fast, adding only up to 1.4 ms overhead to the base QA model.